# Making Reasoning Matter: Measuring and Improving Faithfulness of Chain-of-Thought Reasoning

### v2

@@ 5 Results
We first evaluate the indirect and direct effects of the reasoning chain and reasoning problems on the final outputs. For models (>100B) with the emergent ability to generate plausible reasoning chains, we report natural indirect effects and direct effects (see §2). Table 2 shows the zero-shot performance of the ChatGPT and GPT-4 models. We observe that for StrategyQA and Causal Understanding tasks, GPT-4 has a higher natural indirect effect than ChatGPT, suggesting that it is able to better reason over the reasoning steps for these tasks. However, for mathematical reasoning (GSM8K), ChatGPT has a better indirect effect. Qualitatively, we find that for mathematical reasoning, when we provide intervened reasoning steps, GPT-4 considers them incorrect and continues to generate correct reasoning steps. This results in a lower indirect effect score. Moreover, GPT-4 exhibits a more pronounced direct effect than ChatGPT, suggesting that its outputs are more causally sensitive to reasoning problems.
Table 3 shows the results of causal mediation analysis for 11 different LMs. In these experiments, we examined the causal behaviour using reasoning chains generated by GPT-4 (controlled setting). We find that in-context learning and instruction-tuning improve the indirect effect over models trained only with language modeling objectives (e.g., LLaMA and Mistral), indicating that these methods help the model align better with the reasoning chains. We observe that models trained with RLHF objective (ChatGPT, Llama-2-7B-Chat) have a more direct effect than an indirect effect, suggesting that training on human feedback might have disincentive faithful reasoning (Sharma, 2023). Interestingly, we observe that none of the models has high indirect or direct effects on the causal understanding task. One intuitive reason is that the causal understanding task is challenging, and the model’s (<10B) performance is nearly random; hence, the effects are not strong. Overall, we observe that LLMs are inconsistent in faithfully performing reasoning over the CoT.
@@
Our Model $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
Table 4: Performance of small-sized LMs on four different reasoning tasks. The base models are T5-large (770M) and T5-3B (3B). We report accuracy (${}^{\diamond}$: few-shot performance, $*{}^{*}$: p-value<0.01, ${}^{+}$: p-value<0.05
Comparing Frodo with baselines. We now empirically compare Frodo with three strong baseline models (see Table 4). We consider T5-large (770M) as the inference and reasoning modules. We have the following three observations. First, we present the performance of GPT-$3.5$ on these tasks. We observe the performance on StrategyQA is much lower than on other tasks, indicating the rationales generated for this task can be unfaithful. Hence, similar to (Ramnath et al., 2024), for training Frodo, we use only the instances where the answer predicted by GPT-$3.5$ is correct. Second, for all four datasets, we observe that Frodo outperforms the strong self-rationalization baselines. Frodo, on average, improves the performance by +$4.1$ and +$3$ accuracy points compared to the SFT + CoT and Mario (the strongest baseline), respectively, across all four tasks. Since SFT + CoT and Mario use the same knowledge from GPT-3.5, our results suggest that both our inference and reasoning modules bring substantial performance gains to the model. Third, it is worth noting that increasing (large to 3B) the model size does not hamper the performance of Frodo.
Ablation. To obtain a better insight into the contribution of each component of Frodo, we perform an ablation study (see Table. 4). First, when we do not use the DPO to train our inference module, we see a consistent drop (-$1.9\%$) in performance across the four tasks, indicating the importance of incorporating implicit feedback provided by the DPO in the model’s training. Further, we observe a considerable drop in performance when we do not use counterfactual (-$3.1\%$) and margin ranking loss (-$2.8\%$). This result highlights the model’s ability to benefit from including counterfactual examples.
6 Quantative Analysis
Robustness. In Table 5, we report the controlled indirect effect that indicates how robustly models are able to change their answers when provided with controlled (generated by GPT-4) counterfactual reasoning chains. For StrategyQA, we observe that Frodo significantly improves the robustness performance for T5-3B (+7.7 pp.). Further, for the QuaRel task, we observe +2 pp. improvement over the SFT + CoT method. Qualitatively, we find that for the MCQA tasks, the gold rationales often contain the answer; hence, the SFT + CoT learns to copy those as answers. Further, we perform an ablation to understand which component contributes most to the model’s robustness. We find that counterfactual loss brings the most gain in robustness.
Generalization. The idea is to test our model’s capability to determine if it can improve out-of-distribution (OOD) generalization. Table 6 shows the OOD performance, where we compare our method with SFT+CoT. We trained the models on the OBQA and QASC datasets and evaluated them on the StrategyQA task. We conclude that Frodo significantly helps improve the model’s generalizability to a dataset unseen during fine-tuning.
@@ 6 Quantative Analysis
Faithfulness. Finally, we compare the faithfulness of reasoning chains generated by Frodo with SCoTT, CoT and SFT+COT (see Fig.4). We observed that Frodo achieves a much higher LAS score than the other three baselines, suggesting that DPO training with implicit casual feedback helped the model.
Figure 4: Faithfulness (LAS) of the compared methods on the experiments on StrategyQA datasets. The base Model is T5-3B.
@@ 7 Related Work
Measuring Faithfulness CoT.
Jacovi and Goldberg (2020) argued that obtaining faithful explanations that accurately reflect a model’s reasoning process is important to understand the reasons behind its answer. Recently, Atanasova et al. (2023) proposed a new benchmark to test the faithfulness of natural language explanations. Turpin et al. (2023) proposed identifying examples of unfaithful CoT in adversarially constructed settings, showing that CoT reasoning is not always faithful. To determine faithfulness, they provided bias features in the few-shot setting or made edits to the input. (Lanham et al., 2023) argued that LLM ignores mistakes when introduced into the CoT, which reveals that the LLM is unfaithful. Finally.(Parcalabescu and Frank, 2023) proposed CC-SHAP to measure how well a model’s input contributions align when it produces a prediction and explanation and use it for post-hoc and CoT explanations. Unlike previous work in this paper, we propose measuring the model’s faithful reasoning over the CoT. We propose to use causal mediation analysis to interpret the relationship between the CoT traces and the final answer.
@@
Table 6: Generalization Performance (accuracy) of the compared methods, which are firstly trained on a source dataset and then directly predicting on a target dataset (denoted as source → target).
Self-Rationalization & CoT Distillation.
Initial work on self-rationalization approaches focused on collected gold human rationales and training a model to learn to generate such rationales (Camburu et al., 2018; Wiegreffe et al., 2021; Paul and Frank, 2021). With the advent of LLMs, recently many works have distilled chain-of-thought from LLMs and endowed small LMs with step-by-step reasoning capabilities (Fu et al., 2023; LI et al., 2022; Shridhar et al., 2023; Li et al., 2023). Our work involves distilling CoT from LMs to a smaller one, similar to a certain line of work. We differ in using implicit feedback to enhance the correctness of the distilled CoT.
@@
Causal mediation analysis is an important tool that is used to effectively attribute the causal effect of mediators on an outcome variable (Pearl, 2001). Vig et al. (2020) proposed to use this method to implicate specific neurons and attention heads in mediating gender bias in various pre-trained LMs. Later, this method was used for analyzing different models’ behaviour for different downstream tasks such as Subjective-Verb agreement (Finlayson et al., 2021), Fake News Detection (Chen et al., 2023), arithmetic reasoning (Stolfo et al., 2023), political polarization (Tierney and Volfovsky, 2021). To the best of our knowledge, our study is the first attempt to use casual mediation analysis to analyze the faithfulness of LLMs in their reasoning capabilities.
@@ 8 Conclusion
In this work, we first perform a causal mediation analysis to study the indirect effect of CoT on the final output of twelve LLMs. Our experiments show large variations across tasks and models in how strongly reasoning traces causally affect the model’s prediction. LLMs generally do not reliably use their intermediate reasoning steps when generating an answer. We propose a new method Frodo that tailors small-sized LMs to generate correct reasoning chains and use the reasoning chains as input and faithfully reasons over them to arrive at the correct answer. Experiments show that our method outperforms strong baselines on four reasoning tasks. Our analysis shows that Frodo can outperform the standard supervised knowledge distillation method in out-of-distribution settings.
@@ Acknowledgment
We would like to thank Angelika Romanou, Beatriz Borges, Sahithya Ravi, Gail Weiss, Maxime Peyrard, Syrielle Montariol, Anna Sotnikova, Negar Foroutan and Zeming Chen for their helpful feedback on a draft version of the paper. We acknowledge the support of the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215). West’s lab is partly supported by grants from the Swiss National Science Foundation (200021_185043), Swiss Data Science Center (P22_08), H2020 (952215), Microsoft Swiss Joint Research Center, and Google, and by generous gifts from Facebook, Google, and Microsoft. Antoine Bosselut gratefully acknowledges the support of Innosuisse under PFFS-21-29, the EPFL Science Seed Fund, the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.
@@ 9 Limitations
Compared to training a standard CoT distillation process, our method requires (i) additional counterfactual data generated by LLMs, which can be expensive, and (ii) training time increases as training Direct Preference Optimization is a two-step process. To manage the complexity of our already large-scale experiments involving (a) four different reasoning tasks, and (b) hyperparameter search grids, we ran experiments with 3 random seeds. Additionally, Frodo is dependent on rationales generated by LLMs. Extra care should be taken when applying our model in production environments, especially when making critical decisions or exposing its generated contents directly to human end users.
@@ Appendix A Appendix
A.1 Dataset and Implementation Details

### v4

@@ 5 Results
In Table. 2, 3, we report the results of the causal mediation analysis for twelve models. In section §4, we provide the details about the implementation, evaluation metrics and datasets.
Natural Direct and Indirect effects.
We first evaluate the indirect and direct effects of the reasoning chain and reasoning problems on the final outputs. For models (>100B) with the emergent ability to generate plausible reasoning chains, we report natural indirect effects and direct effects (see §2). Table 2 shows the zero-shot performance of the ChatGPT and GPT-4 models. We observe that for StrategyQA and Causal Understanding tasks, GPT-4 has a higher natural indirect effect than ChatGPT, suggesting that it is able to better reason over the reasoning steps for these tasks. However, for mathematical reasoning (GSM8K), ChatGPT has a better indirect effect. Qualitatively, we find that for mathematical reasoning, when we provide intervened reasoning steps, GPT-4 considers them incorrect and continues to generate correct reasoning steps. This results in a lower indirect effect score. Moreover, GPT-4 exhibits a more pronounced direct effect than ChatGPT, suggesting that its outputs are more causally sensitive to reasoning problems. In general, our experiments show a large variation in the causal effects of COT in the final answer depending on the tasks.
Controlled Direct and Indirect effects.
Table 3 shows the results of causal mediation analysis for 12 different LMs. In these experiments, we examined the causal behaviour using reasoning chains generated by GPT-4 (controlled setting). Our study suggests that vanilla LMs (<20B) (in a zero-shot setting) are systematically unfaithful and consistently fail to reason over the mediator. Increasing the model size (7B to 70B) improves the indirect effect (makes them more faithful), indicating the importance of model size. We find that in-context learning and instruction-tuning improve the indirect effect over models trained only with language modelling objectives (e.g., LLaMA and Mistral), indicating that these methods help the model align better with the reasoning chains. We observe that models trained with RLHF objective (ChatGPT, Llama-2-7B-Chat) have a more direct effect than an indirect effect, suggesting that training on human feedback might have disincentive faithful reasoning (Sharma, 2023). Models that are instruction-tuned or trained on the chain of thought (e.g., Flan-T5) during the pre-training phase have a better indirect effect across different reasoning tasks, suggesting that fine-tuning on CoT can make the model more faithful. Similar to Turpin et al. (2023), we observe inverse scaling for certain tasks. In our case, the indirect effect worsens with increasingly capable models, indicating that sheer scale might not guarantee faithful reasoning. Interestingly, we also observe that none of the models has high indirect or direct effects on the causal understanding task. One intuitive reason is that the causal understanding task is challenging, and the model’s (<10B) performance is nearly random; hence, the effects are not strong. Overall, we observe that LLMs are inconsistent in faithfully performing reasoning over the CoT.
@@
Frodo 82.1 93.5 80.1 75.9
LlaMa-2-7B 67.2 56.8 47.5 49.6
SFT + CoT 79.4 68.4 62.8 54.6
Frodo 81.5+ 73.5+ 71.4+ 63.4+
Mistral-7B 58.2 56.8 82.1 65.2
SFT + CoT 78.2 70.8 83.5 70.1
Frodo 81.9+ 78.2+ 84.9+ 72.3+
Table 4: Performance of small-sized LMs (770M-7B) on four different reasoning tasks. The base models are T5-large (770M), T5-3B (3B), LLaMa-2-7B and Mistral-7B . We report accuracy (%).: few-shot performance, : p-value<0.01, +: p-value<0.05
Comparing Frodo with Baselines. We now empirically compare Frodo with three strong baseline models (see Table 4). We consider T5-large (770M) as the inference and reasoning modules. We have the following three observations. First, we present the performance of GPT-$3.5$ on these tasks. We observe the performance on StrategyQA is much lower than on other tasks, indicating the rationales generated for this task can be unfaithful. Hence, similar to (Ramnath et al., 2024), for training Frodo, we use only the instances where the answer predicted by GPT-$3.5$ is correct. Second, for all four datasets, we observe that Frodo outperforms the strong self-rationalization baselines. Frodo, on average, improves the performance by +$4.1$ and +$3$ accuracy points compared to the SFT + CoT and Mario (the strongest baseline), respectively, across all four tasks. Since SFT + CoT and Mario use the same knowledge from GPT-3.5, our results suggest that both our inference and reasoning modules bring substantial performance gains to the model. Third, it is worth noting that increasing (770M to 3B) the model size does not hamper the performance of Frodo. Fourth, we also report the performance of the LLaMa-2-7B and Mistal-7B models. We show that FRODO further improves the performance of model size 7B.
@@
Ablation.
To obtain a better insight into the contribution of each component of Frodo, we perform an ablation study (see Table. 4). First, when we do not use the DPO to train our inference module, we see a consistent drop (-$1.9\%$) in performance across the four tasks, indicating the importance of incorporating implicit feedback provided by the DPO in the model’s training. Further, we observe a considerable drop in performance when we do not use counterfactual (-$3.1\%$) and margin ranking loss (-$2.8\%$). This result highlights the model’s ability to benefit from including counterfactual examples.
6 Analysis
6.1 Quantitative Analysis
Robustness.
In Table 5, we report the controlled indirect effect that indicates how robustly models are able to change their answers when provided with controlled (generated by GPT-4) counterfactual reasoning chains. For StrategyQA, we observe that Frodo significantly improves the robustness performance for T5-3B (+7.7 pp.). Further, for the QuaRel task, we observe +2 pp. improvement over the SFT + CoT method. Qualitatively, we find that for the MCQA tasks, the gold rationales often contain the answer; hence, the SFT + CoT learns to copy those as answers. Further, we perform an ablation to understand which component contributes most to the model’s robustness. We find that counterfactual loss brings the most gain in robustness.
Generalization.
The idea is to test our model’s capability to determine if it can improve out-of-distribution (OOD) generalization. Table 7 shows the OOD performance, where we compare our method with SFT+CoT. We trained the models on the OBQA and QASC datasets and evaluated them on the StrategyQA task. We conclude that Frodo significantly helps improve the model’s generalizability to a dataset unseen during fine-tuning.
6.2 Qualitative Analysis
Causal Analysis.
To understand the reason for the inconsistency in the causal effect, we analyze its relationship with problem complexity. In Table 6, we report the indirect effect of CoT with respect to the number of reasoning steps for GSM8K problems. We observe that with the increase in the number of reasoning steps, the indirect effect drops for both LLaMa-3 and ChatGPT. It indicates that the length of the reasoning steps has an inverse effect on the faithfulness of these models. Table 16 shows a few examples of different models’ unfaithful reasoning over the chain of thought.
No. RS LLaMa-3-70B ChatGPT
2 66.92 65.6
3 50.31 53.9
4 43.2 61.2
5 41.8 55
6 37.5 21
7 0 25
overall 57.2 56.3
Table 6: Indirect Effect. No. RS = Number of Reasoning Steps. IE: $Avg[Acc(Y_{00})-Acc(Y_{01})]$
FRODO Analysis. To further understand the findings in §5, we manually analyze the relevance of the $100$ CoT generated by SFT and DPO. We observed that SFT generates 74% and 54% relevant CoT, whereas DPO generated CoT 77% and 59% relevant for QuaRel and OBQA tasks, respectively. Further, we observe two types of errors made by SFT and DPO: (i) invalid reasoning steps– reasoning steps leading to incorrect answers and (ii) unnecessary steps – reasoning steps not containing enough information to support a correct answer. For DPO, we observed that 40% of the errors are invalid reasoning steps, and 56% are unnecessary steps. Table 21 shows some examples of CoT generated by SFT and DPO, SFT generated CoT are incomplete or contradictory.
@@ 7 Related Work
Measuring Faithfulness CoT. Jacovi and Goldberg (2020) argued that obtaining faithful explanations that accurately reflect a model’s reasoning process is important to understand the reasons behind its answer.(Atanasova et al., 2023) proposed a new benchmark to test the faithfulness of natural language explanations. Turpin et al. (2023) proposed identifying examples of unfaithful CoT in adversarial settings, showing that CoT reasoning is not always faithful. To determine faithfulness, they provided bias features in the few-shot setting or made edits to the input. (Lanham et al., 2023) argued that LLM ignores mistakes when introduced into the CoT, which reveals that the LLM is unfaithful. Finally, (Parcalabescu and Frank, 2024) introduced CC-SHAP to measure input alignment with predictions for both post-hoc and CoT explanations. Unlike prior work, we employ causal mediation analysis to measure the model’s faithful reasoning over the CoT, and to interpret its relationship with the answer.
@@
Table 7: Generalization Performance (accuracy) of methods, trained on a source dataset and directly predicting on a target dataset (denoted as source → target).
Self-Rationalization and CoT Distillation.
Initial work on self-rationalization approaches focused on collected gold human rationales and training a model to learn to generate such rationales (Wiegreffe et al., 2021; Paul and Frank, 2021; Camburu et al., 2018). With the advent of LLMs, recently many works have distilled CoT from LLMs and endowed small LMs with step-by-step reasoning capabilities (Fu et al., 2023; LI et al., 2022; Shridhar et al., 2023; Li et al., 2023). Our work involves distilling CoT from LMs to a smaller one, similar to a certain line of work. We differ in using implicit feedback to enhance the correctness of the distilled CoT.
@@
Causal mediation analysis is an important tool that is used to effectively attribute the causal effect of mediators on an outcome variable (Pearl, 2001). Vig et al. (2020) proposed to use this method to implicate specific neurons and attention heads in mediating gender bias in various pre-trained LMs. Later, this method was used for analyzing different models’ behaviour for different downstream tasks such as Subjective-Verb agreement (Finlayson et al., 2021), Fake News Detection (Chen et al., 2023), arithmetic reasoning (Stolfo et al., 2023), political polarization (Tierney and Volfovsky, 2021). To the best of our knowledge, our study is the first attempt to use casual mediation analysis to analyze the faithfulness of LLMs in their reasoning capabilities. In this work, we followed Pearl (2001) to perform the mediation analysis. The mediation analysis allows us to measure the following: Direct effect: Contribution of X (input) to Y (output). Indirect effect: Contribution of R (reasoning chain) to Y (output). Hence, a high direct effect means the model’s output is primarily influenced by the input, and a high indirect effect means the reasoning chain has more effect on the output.
@@ 8 Conclusion
In this work, we perform a causal mediation analysis to study the indirect effect of CoT on the final output of twelve LLMs. Our experiments show large variations across tasks and models in how strongly reasoning traces causally affect the model’s prediction. LLMs generally do not reliably use their intermediate reasoning steps when generating an answer. We introduce Frodo that tailors small-sized LMs to generate correct reasoning chains and faithfully reason over them to arrive at the correct answer. Experiments show that our method outperforms strong baselines on four reasoning tasks, including out-of-distribution settings.
@@ Acknowledgment
We would like to thank Angelika Romanou, Beatriz Borges, Sahithya Ravi, Gail Weiss, Maxime Peyrard, Syrielle Montariol, Anna Sotnikova, Negar Foroutan and Zeming Chen for their helpful feedback on a draft version of the paper. We acknowledge the support of the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215). West’s lab is partly supported by grants from the Swiss National Science Foundation (200021_185043), Swiss Data Science Center (P22_08), H2020 (952215), Microsoft Swiss Joint Research Center, and Google, and by generous gifts from Facebook, Google, and Microsoft. Antoine Bosselut gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.
@@ 9 Limitations
A limitation of our Causal Analysis metric is that it does not evaluate the model’s real internal reasoning process. Without a complete understanding of the pertaining data and models’ internal working process, it is difficult to know whether or not the chain of thought is faithful to the reasoning process. In this study, we provide insist and evidence that could explain how the model uses CoT. For future research, the causal mediation metric can be useful for measuring the extent to which new methods improve faithfulness. Compared to training a standard CoT distillation process, our method requires (i) additional counterfactual data generated by LLMs, which can be expensive, and (ii) training time increases as training Direct Preference Optimization is a two-step process. To manage the complexity of our already large-scale experiments involving (a) four different reasoning tasks, and (b) hyperparameter search grids, we ran experiments with 3 random seeds. Additionally, Frodo is dependent on rationales generated by LLMs. Extra care should be taken when applying our model in production environments, especially when making critical decisions or exposing its generated contents directly to human end users.
@@ Appendix A Appendix
A.1 FRODO - Inference Module (DPO)
The preference data of human or artificial annotators is modeled by a learnable implicit reward model $f_{\theta}$ under Bradley-Terry theories (Bradley and Terry, 1952):
A.2 Additional Experimental Results
RQ1: How faithful FRODO is compared to SCoTT and CoT?
Finally, we compare the faithfulness of reasoning chains generated by Frodo with SCoTT, CoT and SFT+COT (see Fig.4). We observed that Frodo achieves a much higher LAS score than the other three baselines, suggesting that DPO training with implicit casual feedback helped the model.
RQ2: How does FRODO work on GSM8k?
Table 8 reports the performance of FRODO on math reasoning problems. We observe that FRODO outperform SFT by +$3.75$ (average).
Model GSM8K
LLama-2 7B + SFT + COT 17.8
LLama-2 7B + SFT + FRODO 21.1
Mistral + SFT + COT 40.4
Mistral + SFT + FRODO 44.6
Table 8: Performance of FRODO on GSM8K (accuracy)
RQ3: How does FRODO generalize on Causal Understanding Task?
Table 9 reports the result of FRODO on the Causal Understanding task. Please note that the Causal Understanding dataset does not have training data. Hence, we trained FRODO on the StrategyQA dataset and evaluated it on Causal Understanding. Our results suggest that FRODO can generalize better than SFT.
Model SQ → CU
SFT + COT 51.0
SFT + FRODO 53.2
Table 9: Generalization Performance. (accuracy) of methods, trained on a source dataset and directly predicting on a target dataset (denoted as source → target), where SQ = StrategyQA and CU = Causal Understanding.
RQ4: How well did GPT-4 generate the Chain-of-Thought? We manually evaluated the quality of the chain of thought generated by GPT-4. We found 94% correct CoTs generated by GPT-4 for StrategyQA, whereas only 65% correct for the Causal Understanding task. This also indicates why the performance of GPT-4 drops for the Causal Understanding task (see Table 2).
SQ GSM8K CU
94% 82% 65%
Table 10: Human Judgement of GPT-4 reasoning chain. SQ: StrategyQA, CU: Causal Understanding.
RQ5: What are the differences between FRODO and Selection-Inference (Creswell et al., 2023) method?
The key differences between FRODO and the Selection-Inference method are:
  1. 1.
    The selection-inference framework assumes that each question is accompanied by context information, which contains all the information necessary to solve the problem.
  2. 2.
    FRODO does not have that assumption; therefore, our method works on open-domain question-answer tasks. Hence, we compare our method with RAINER, CRYSTAL, and MARIO, which do not have such assumptions.
In the selection-inference framework, SFT with language modelling loss is used to train the inference module, while we used counterfactual loss, LM loss, and margin-ranking loss.
RQ6: How is the performance of FRODO on the Entailmentbank dataset?
Table 11 shows that FRODO clearly outperforms previous baselines on the Entailmentbank dataset.
Models Task 1 Task 2
Entailment Writer 34.4 23.2
METGEN 37.0 28.0
FRODO 38.8 34.5
Table 11: Performance of FRODO on Entailment Bank dataset.
Figure 4: Faithfulness (LAS) of the compared methods on StrategyQA. The base Model is T5-3B.
RQ7: What is the causal effect of open-sourced models (<100B) in natural setting?
We experimented with open-sourced larger models LLaMa-3-70B (zero-shot setting) for natural indirect effect. However, we have observed that 70B models do not have the emergent abilities to generate coherent/meaningful reasoning steps in zero-shot settings. Similar to our observation, previous studies (Kojima et al., 2022) also discussed that only models >100B start showing such emergent abilities. Table 12 reports the Natural Indirect effect of LLaMA-2 and LLaMA-3 (70B).
Models Causal Effects StrategyQA
LLaMa-3 70B NIE 21.1
LLaMa-3 70B CIE 23.1
LLaMa-2 70B NIE 12.1
LLaMa-2 70B CIE 24.1
Table 12: Performance of FRODO on GSM8K (accuracy)
A.3 Dataset and Implementation Details